# Discovery of a dwarf planet candidate in an extremely wide orbit: 2017 OF201

### v1

@@ Abstract
We report the discovery of a dwarf planet candidate, 2017 OF201, currently located at a distance of 90.5 au. Its orbit is extremely wide and extends to the inner Oort cloud, with a semi-major axis of 838 au and a perihelion of 44.9 au precisely determined from 19 observations over seven years. Assuming a typical albedo of 0.15, we estimate a diameter $\sim$700 km, making it the second-largest known object in this dynamical population and a likely dwarf planet. Its high eccentricity suggests that it is part of a broader, unseen population of similar objects totaling $\sim$1of Earth’s mass. Notably, the longitude of perihelion of 2017 OF201 lies well outside the clustering observed in extreme trans-Neptunian objects, which has been proposed as dynamical evidence for a distant, undetected planet.
@@
facilities: Blanco (DECam), CFHT (MegaCam)software: astropy (Astropy Collaboration et al., 2013), matplotlib (Hunter, 2007), scipy (Jones et al., 2001), SExtractor (Bertin & Arnouts, 1996), PSFEx (Bertin, 2011), SCAMP (Bertin, 2006), astrometry.net (Lang et al., 2010).
{CJK*}
UTF8gkai
@@
show]qy8274@princeton.edu
1 Introduction
@@
While the asteroid belt contains only $\sim$0.04of the Earth’s mass ($M_{\oplus}$), the Kuiper Belt – home to Pluto and numerous icy bodies – harbors significantly more material, with an estimated total mass of 2% $M_{\oplus}$ (Fraser et al., 2014; Pitjeva & Pitjev, 2018). Even greater mass is believed to reside beyond the classic Kuiper Belt, in the scattering disk and Oort Cloud, which may together contain up to several $M_{\oplus}$ (Gladman & Volk, 2021). These distant reservoirs of comets, small bodies, minor planets and possibly undetected massive planets holds essential clues to the formation and evolution of the solar system.
@@
After 30 years of the discovery of the first trans-Neptunian objects (TNOs) (Jewitt & Luu, 1993) other than Pluto, various survey programs have been conducted to explore the vast region of outer solar system, and more than 5,000 TNOs have been discovered so far. Although the census of large TNOs near the ecliptic plane is nearly complete, it remains highly incomplete in regions that are beyond $\sim$ 60 au or at high latitude, due to the challenging scaling relation that the brightness of reflected light drops steeply with distance as $r^{-4}$. Many dedicated solar-system surveys are thus either shallower wide-field surveys covering $\sim$10,000 deg2 (e.g., Trujillo & Brown, 2003; Schwamb et al., 2010) or deeper but more focused surveys near the ecliptic plane or pencil-beam regions covering hundreds to a few thousand deg2 (e.g., Elliot et al., 2005; Petit et al., 2011; Bannister et al., 2018; Sheppard & Trujillo, 2016; Sheppard et al., 2019).
@@
Fortunately, recent cosmological imaging programs, though not dedicated to solar-system science, offer a promising solution to this limitation. These surveys typically provide wide sky coverage, deep photometry, and multiple epochs that are well-suited for the detection of faint solar system objects. A prime example is the Dark Energy Survey (Dark Energy Survey Collaboration et al., 2016), whose main scientific goal lies in weak lensing cosmology but has already led to the discovery of $\sim 800$ TNOs (Bernardinelli et al., 2022).
In this letter, we report the discovery of a large and exotic TNO, 2017 OF201  [^1] https://www.minorplanetcenter.net/mpec/K25/K25K47.html, from a systematic search in the archival data of another extra-galactic survey similar to DES. This survey is called the Dark Energy Camera Legacy Survey [^2] https://www.legacysurvey.org/ (DECaLS; Dey et al., 2019), which is a wide-field optical imaging survey covering mostly $-15^{\circ}<\mathrm{Dec}<+32^{\circ}$ in the $grz$-bands using the Dark Energy Camera (DECam) on the 4-meter Blanco telescope (Flaugher et al., 2015). To our knowledge, this dataset has not been previously searched for solar-system objects, possibly due to its sparse and irregular epoch sampling.
We shall describe the discovery of 2017 OF201 in Section 2, present its large size and wide orbit in Section 3, explore its possible migration history in Section 4, and finally discuss its interesting implications for the Planet X / Planet 9 hypothesis.
2 Discovery
@@
Figure 1: Trajectory of 2017 OF201 on the sky from 2011 to 2018. Individual detections from 13 nights are shown on top of the predicted trajectory based on the best-fit orbit, which describes the detections very well with a scatter of 0.13 (DECam) and 0.03 (CFHT) arcsec in each component, consistent with the estimated astrometric error. The insets show example images from DECam ($r$-band on 2017-09-17) and CFHT ($r$-band on 2011-08-31).
@@
We performed our search in a similar way to Bernardinelli et al. (2020, 2022) and optimized it for the sparse epoch sampling of DECaLS. Details of the search algorithm will be described in a future paper. So far, 2017 OF201 is the most interesting object found in our search, with the largest distance and widest orbit.
@@
After identifying ten detections of 2017 OF201 in three bands ($g,r,z$) of DECaLS span from 2014 to 2018 and connecting them with a tentative orbit, it became rather clear that they correspond to a single moving object that has an extremely wide and eccentric orbit. In addition, with an apparent magnitude $r\sim$ 22.6 mag at a distance of 85 au at the observation epochs, 2017 OF201 has the second brightest absolute magnitude among 411 known TNOs with wide orbits ($a>$ 80 au) in the JPL Small-Body Database [^3] https://ssd.jpl.nasa.gov/tools/sbdb_query.html.
@@
With the ephemeris derived from DECam detections, we were able to quickly search for additional detections in the data archives of other telescopes to further improve the precision of the orbital fit and photometric measurements. We found nine additional $r$-band images from the data archive of the 3.6 m Canada-France-Hawaii Telescope (CFHT), in which 2017 OF201 is found exactly at the predicted positions. These images were taken in 2011 and 2012, increasing the observation arc to seven years. We have also searched the data archive of the Subaru telescope and Gemini-North telescope, but unfortunately found no coverage.
@@
$a$ (au) 838.3 $\pm$ 6.8
$e$ 0.94643 $\pm$ 0.00045
$i$ (deg) 16.20509 $\pm$ 0.00009
$\Omega$ (deg) 328.59148 $\pm$ 0.0007
$\omega$ (deg) 337.73091 $\pm$ 0.0016
$\varpi$ (deg) 306.32238 $\pm$ 0.0023
perihelion $q$ (au) 44.9093 $\pm$ 0.0086
aphelion $Q$ (au) 1632 $\pm$ 14
period (yrs) 24256
arc length (days) 2618
reference epoch $t_{0}$ (MJD) 56932.524451
(2014-10-02)
distance at $t_{0}$ (au) 83.83309 $\pm$ 0.00049
radial velocity at $t_{0}$ (au/yr) 0.6355 $\pm$ 0.0002
$V$ mag at $t_{0}$aaComputed using the Find_Orb software. 22.8
$H_{V}$aaComputed using the Find_Orb software. 3.5
$g-r$ 0.77 $\pm$ 0.11
$r-z$ 0.80 $\pm$ 0.07
$B-V$bbfootnotemark: 0.99 $\pm$ 0.11
aafootnotetext: Converted from $g-r$ using the SDSS transformation formula from Smith et al. (2002).
3 Properties of 2017 OF201
We performed precise astrometric and photometric measurements (see Appendix A) on each of the 19 detection images to determine the orbit, color, and absolute magnitude of 2017 OF201. A summary of the derived properties is listed in Table 1.
Figure 1 illustrates the trajectory of 2017 OF201 on the sky, marked with detection dates and example detection images from DECam and CFHT. The combination of motion patterns from parallax (elliptic) and proper motion (straight) is recognizable. In fact, 2017 OF201 is quickly moving away from us and its parallax shrank by 5% over the 7-year observation arc, enabling a precise determination of radial velocity and therefore the full set of orbital parameters.
@@ 3 Properties of 2017 OF201
The last time 2017 OF201 passed close to us was in November of 1930, and it will come back again in about 25,000 years. Interestingly, this last perihelion was in the same year of Pluto’s discovery (Feb 1930), when 2017 OF201 reached its maximum brightness of $V=20.1$ mag, about 4 magnitudes fainter than Pluto [^4] Coincidentally, it was also in the same year of the establishment of the Institute for Advanced Study, where the first author of this paper is based..
3.1 Orbit
@@
The orbit of 2017 OF201 is determined using the online version of the Find_Orb software [^5] https://www.projectpluto.com/fo.htm, which implements a least-squares fitting of the orbital elements to the astrometric measurements of the 19 detections. The best-fit values and uncertainty of orbital elements are listed in Table 1. Thanks to the long observation arc from 2011 to 2018, the orbit of 2017 OF201 is precisely determined, with an undoubtedly wide semi-major axis ($a=838\pm 7$ au), moderate perihelion distance ($q=44.9$ au), and low inclination ($i=16.2^{\circ}$), located at the boundary between the scattering disk and inner Oort cloud.
@@
The left panel of Figure 2 shows the plan view projection of the orbit of 2017 OF201 onto the ecliptic plane, together with some other TNOs with extremely wide and eccentric orbits (termed “extreme TNOs”) for comparison. Its orbit is most similar to that of 2013 SY99 in terms of the semi-major axis and eccentricity (Bannister et al., 2017). With such a large aphelion distance ($\sim 1600$ au), the Galactic tide will start to affect the orbital evolution. We will discuss its orbital dynamics in Section 4.
@@
A notable characteristic of the orbit of 2017 OF201 is its longitude of perihelion of $\varpi=306^{\circ}$, which significantly deviates from the apparent clustering observed among other extreme TNOs first noticed by Trujillo & Sheppard (2014). The right panel of Figure 2 illustrates the distribution of longitude of perihelia $\varpi$ and semi-major axes $a$ for known extreme TNOs. Objects with $a>200$ au are predominantly clustered around $\varpi\approx 60^{\circ}$, which has been suggested as evidence for an undetected super-Earth beyond several hundred au (e.g., Trujillo & Sheppard, 2014; Brown & Batygin, 2016; Batygin & Brown, 2016; Sheppard et al., 2019; Siraj et al., 2025). We shall explore later in Section 4.2 the implications of the fact that 2017 OF201 is an outlier to this clustering.
@@
Figure 2: Left: Plan view of the orbits of TNOs with extremely wide orbits, including our newly discovered 2017 OF201, which has a distinct orbit is an outlier to the apsidal clustering of the others. For reference, the most probable orbit of Planet X from Siraj et al. (2025) is shown in black. Right: Distribution of longitude of perihelion $\varpi$ of TNOs from Siraj et al. (2025). The purple shaded region indicates the clustering of $\varpi$ around 60. Objects with (meta)stable orbits are shown as solid (open) symbols. The newly discovered 2017 OF201, highlighted in red, does not belong to the apsidal clustering.
3.2 Color and variability
@@
2017 OF201 has a color index of $g-r=0.77\pm 0.11$ mag, similar to Sedna ($g-r=0.85$) and much redder than the Sun ($g-r=0.44$; Willmer, 2018). Compared with the color distribution of the TNO population in Sheppard (2010), 2017 OF201 has a color consistent with other TNOs in the scattering disk and detached orbits, but is on the redder side. Given the depth and time sampling of the DECam and CFHT data, we do not detect any variability over 0.1 mag, which means the shape of 2017 OF201 is likely close to spherical. Deeper imaging with higher cadence is needed to probe the short-time variability due to rotation or satellite transit.
3.3 Size and mass estimate
@@
The discovery of such a large object with a high orbital eccentricity has interesting implications for the population of objects in the trans-Neptunian space. Throughout the orbit of 2017 OF201, it is detectable in DECaLS for only $\sim$0.5of the period around perihelion. This limited visibility window strongly suggests that a substantial population of similar objects – with large sizes, wide orbits, and high eccentricities – should exist but be difficult to detect due to their extremely large distance.
The mass of 2017 OF201 is estimated to be 3$\times$1020 kg = 1/20,000 $M_{\oplus}$ when assuming a typical density of $1.7\ \mathrm{g\,cm^{-3}}$ and a diameter of 700 km. Therefore, the whole population behind 2017 OF201 would add up 200 times to 1% $M_{\oplus}$ (about the Moon’s mass). For comparison, the total mass of the classic Kuiper belt is estimated to be 1–2% $M_{\oplus}$ from object counting (Bernstein et al., 2004; Fraser et al., 2014) and dynamical estimates (Pitjeva & Pitjev, 2018), and that of the scattering disk is 1–10% $M_{\oplus}$ (Gladman et al., 2008; Volk & Malhotra, 2008). Discovery of large TNOs with extremely elongated orbits such as 2017 OF201 and Sedna (Brown et al., 2004) thus reveals significant amount of mass in the outer solar system.
4 Orbital dynamics
We investigate the long-term dynamical evolution of 2017 OF201 using N-body simulations with the ias15 integrator (Rein & Spiegel, 2015) in REBOUND [^9] https://rebound.readthedocs.io/ (Rein & Liu, 2012). Only the outer two giant planets (Uranus, Neptune) are included explicitly, while the combined influence of the inner giants (Jupiter, Saturn) is modeled as a static solar oblateness ($J_{2}$), following Batygin & Brown (2016) and Gerdes et al. (2017) [^10] For validation purposes, we also run computationally intensive simulations that directly integrated 2017 OF201 with Jupiter, Saturn, Uranus, Neptune, and Galactic tide. These simulations yielded results that are qualitatively consistent with our simplified model.. Due to 2017 OF201’s large aphelion distance ($\sim 1600$ au), its orbit may be perturbed by the Galactic tide. We implement the Galactic tidal model from Levison et al. (2001) using the add_force module in REBOUNDx [^11] https://reboundx.readthedocs.io/ (Tamayo et al., 2020).
4.1 Possible origin of 2017 OF201
The semi-major axis of 2017 OF201 can evolve through chaotic diffusion driven by interactions with Neptune. However, the diffusion timescale $T$ increases steeply with perihelion distance. Following Hadden & Tremaine (2024), it is approximately

### v2

@@ Abstract
We report the discovery of a dwarf planet candidate, 2017 OF201, currently located at a distance of 90 au. Its orbit is extremely wide and extends to the inner Oort cloud, with a semi-major axis of 830 au and a perihelion of 45 au, precisely determined from 24 observations over 20 years. Assuming a typical albedo of 0.13, we estimate a diameter $\sim$700 km, making it the second-largest known object in this dynamical population and a dwarf planet candidate with the widest orbit. Its high eccentricity suggests that an unseen population of similar objects would total $\sim$1% of Earth’s mass. Notably, the longitude of perihelion of 2017 OF201 lies outside the clustering observed in extreme trans-Neptunian objects, posing a challenge to the proposed dynamical evidence for the hypothetical Planet Nine.
@@
journal: ApJLfacilities: Blanco (DECam), CFHT (MegaCam), Magellansoftware: astropy (Astropy Collaboration et al., 2013), matplotlib (Hunter, 2007), scipy (Jones et al., 2001), SExtractor (Bertin and Arnouts, 1996), PSFEx (Bertin, 2011), SCAMP (Bertin, 2006), astrometry.net (Lang et al., 2010).
@@
show]eritas.yang@princeton.edu
I Introduction
@@
While the asteroid belt contains only less than 10-3 of Earth’s mass ($M_{\oplus}$), the Kuiper Belt harbors fifty times more material (Pitjeva and Pitjev, 2018). In the vast region beyond, even greater mass is believed to reside in the scattering disk and Oort Cloud, likely totaling several $M_{\oplus}$ (Gladman and Volk, 2021). These distant reservoirs of comets, minor planets, dwarf planets, and possibly undetected massive planets hold essential clues to the formation and evolution history of the solar system. Three decades ago, the first trans-Neptunian objects (TNOs) beyond Pluto was discovered (Jewitt and Luu, 1993). Since then, thousands more have been identified [^1] https://www.minorplanetcenter.net/iau/lists/TNOs.html through various surveys, revealing a remarkable richness of this population.
@@
However, the census of TNOs is far from complete in regions beyond about 60 au or at high ecliptic latitudes, even for large objects such as those qualifying as dwarf planets. Due to the challenging scaling relation that the brightness of reflected light drops steeply with distance as $r^{-4}$, dedicated solar-system surveys are usually either wide (covering $\sim$10,000 deg2) but shallower (e.g., Trujillo and Brown, 2003; Schwamb et al., 2010), or deeper but more focused on the ecliptic plane (covering hundreds to a few thousand deg2, e.g., Elliot et al., 2005; Petit et al., 2011; Bannister et al., 2018; Sheppard and Trujillo, 2016; Sheppard et al., 2019).
@@
Fortunately, recent cosmological imaging surveys, though not dedicated to solar-system science, offer a promising solution to this limitation. With wide sky coverage, deep photometry, and multiple epochs, those surveys are well-suited for detections of faint solar system objects. A prime example is the Dark Energy Survey (DES; Dark Energy Survey Collaboration et al., 2016), whose main scientific goal lies in weak lensing cosmology but has led to the discovery of more than $800$ TNOs (Bernardinelli et al., 2022).
In this letter, we report the discovery of a large TNO, 2017 OF201  [^2] https://www.minorplanetcenter.net/mpec/K25/K25K47.html, out of a systematic search in another extra-galactic survey similar to DES but not searched before. This object is likely large enough to qualify as a dwarf planet, and its extremely wide orbit extends to the inner Oort cloud. We also discuss its migration history and interesting implications for the Planet Nine / Planet X hypothesis (Batygin and Brown, 2016).
@@
Figure 1: Trajectory of 2017 OF201 on the sky from 2011 to 2025, which is a combination of parallax (elliptic) and proper motion (straight) components. Individual detections from 16 nights are shown with the predicted trajectory based on the best-fit orbit, with a small residue of 0.13 (DECam) and 0.03 (CFHT) arcsec in each component, consistent with their astrometric error. The insets show example images from Magellan, DECam, and CFHT.
II Discovery
We performed a systematic search for moving objects in the archival data of the Dark Energy Camera Legacy Survey (DECaLS; Dey et al., 2019) [^3] https://www.legacysurvey.org/decamls/. It uses the Dark Energy Camera (DECam) on the 4-meter Blanco telescope (Flaugher et al., 2015), the same facility as DES, and its sky coverage is about two times larger than DES (declination between $15^{\circ}$ and $32^{\circ}$). This dataset has not been searched for TNOs with its full potential, due to its sparse and irregular epoch sampling, which may separate by months to years.
@@ II Discovery
Our search algorithm is similar to Bernardinelli et al. (2020b, 2022), but optimized for the sparse epoch sampling of DECaLS. Details of the algorithm will be described in a future paper. So far, 2017 OF201 is the most interesting object found in our search, with the largest distance and widest orbit.
@@
After identifying 10 detections in three bands ($g,r,z$) in DECam data between 2014 and 2018, it became clear that they correspond to a single moving object with an extremely wide orbit. With an apparent magnitude $r\sim$ 22.6 mag at 85 au away, 2017 OF201 has the second brightest absolute magnitude among 442 known TNOs with wide orbits ($a>$ 80 au) in the JPL Small-Body Database [^4] https://ssd.jpl.nasa.gov/tools/sbdb_query.html.
@@
To further improve the orbital and photometric fitting, we used the ephemeris derived from the DECam detections to quickly search in other data archives. Following suggestions from Mike Alexandersen through a private communication, we found 9 additional $r$-band images taken in 2011 and 2012 with the 3.6 m Canada–France–Hawaii Telescope (CFHT), where 2017 OF201 is located exactly at the predicted positions [^5] Taken for a TNO project by Alexandersen et al. (2016). It is a pity that their team missed the discovery of 2017 OF201.. The search was performed using the Solar System Object Image Search (SSOIS) tool [^6] https://www.cadc.hia.nrc.gc.ca/en/ssois/ (Gwyn et al., 2012). We also searched the data archive of the Subaru and Gemini-North telescope but unfortunately found no coverage.
After our announcement of 2017 OF201, an amateur astronomer, Sam Deen, found two detections in the Sloan Digital Sky Survey (SDSS; York et al., 2000) [^7] http://www.sdss.org/, taken as early as 2004 and 2009.888In SDSS’s object catalog, these two detections have object IDs 1237666274738438899 and 1237680073395798979, misclassified as a star and a galaxy, respectively. Together with our follow-up observations taken with the 6.5 m Magellan telescope on July 25, 2025, the observational arc has accumulated to 20 years with 24 observations. This unusually long baseline allows us to precisely determine its orbit right after the discovery.
III Properties of 2017 OF201
@@
$a$ (au) 830.1 $\pm$ 0.8
$e$ 0.94589 $\pm$ 0.00005
$i$ (deg) 16.20508 $\pm$ 0.00006
$\Omega$ (deg) 328.5914 $\pm$ 0.0004
$\omega$ (deg) 337.7290 $\pm$ 0.0012
$\varpi$ (deg) 306.3204 $\pm$ 0.0013
perihelion $q$ (au) 44.9189 $\pm$ 0.0012
aphelion $Q$ (au) 1615.2 $\pm$ 1.5
period (yrs) 23900 $\pm$ 30
arc length 7612 days (20.84 years)
reference epoch $t_{0}$ (MJD) 60881.5 (2025-07-25)
mean anomaly $M$ at $t_{0}$ 1.4271 $\pm$ 0.0019
distance at $t_{0}$ (au) 90.7
radial velocity at $t_{0}$ (au/yr) 0.64
$V$ mag at $t_{0}$ 23.3
$H_{g}$ 4.10 $\pm$ 0.09
$H_{r}$ 3.33 $\pm$ 0.05
$H_{i}$ 2.77 $\pm$ 0.10
$H_{z}$ 2.47 $\pm$ 0.10
$H_{V}$ 3.72 $\pm$ 0.09
grg-r 0.77 $\pm$ 0.08
rir-i 0.56 $\pm$ 0.07
BVB-V 0.96 $\pm$ 0.09
Table 1 gives a summary of properties of 2017 OF201 derived from astrometric and photometric measurements of the 22 detection images (Appendix A) and information read directly from SDSS’s catalog.
Figure 1 shows the detections and best-fit orbit of 2017 OF201. It is moving quickly away from us, resulting in a parallax shrink by 15% over 20 years. This significant parallax evolution enables a precise determination of radial velocity – the most uncertain element to measure for TNOs.
@@
The last time 2017 OF201 passed close to us was in November of 1930, and it will come back again in about 25,000 years. Interestingly, this last perihelion was in the same year of Pluto’s discovery (Feb 1930), when 2017 OF201 reached its maximum brightness of $V=20.2$ mag, about 4 magnitudes fainter than Pluto. Coincidentally, it was also in the same year of the establishment of the Institute for Advanced Study, where the first author of this paper is based.
III.1 Orbit
@@
The orbit of 2017 OF201 is determined using the online version of the Find_Orb software [^9] https://www.projectpluto.com/fo.htm, which implements a least-squares fitting of the orbital elements to the astrometric measurements of the 24 detections. Thanks to the very long observation arc from 2004 to 2025, the orbit of 2017 OF201 is precisely determined, with the semi-major axis determined to 0.1% precision ($a=830\pm 1$ au), revealing an undoubtedly wide orbit with moderate perihelion distance ($q=44.9$ au) and low inclination ($i=16.2^{\circ}$).
@@
Figure 2 shows a comparison between orbits of 2017 OF201 and some other TNOs with extremely wide and eccentric orbits (termed “extreme TNOs”). With such a large aphelion distance ($\sim 1600$ au), the Galactic tide may affect the orbital evolution, placing 2017 OF201 at the boundary between the scattering disk and inner Oort cloud. We shall discuss its dynamical history in Section IV.1.
@@
Another notable characteristic of 2017 OF201’s orbit is its longitude of perihelion $\varpi=306^{\circ}$, an outlier to the apparent clustering around $\varpi\approx 60^{\circ}$ claimed among other extreme TNOs (Figure 2, right panel). First noticed by Trujillo and Sheppard (2014), the clustering has been suggested as evidence for an undetected Planet Nine (also known as Planet X) at several hundred au (e.g., Trujillo and Sheppard, 2014; Brown and Batygin, 2016; Batygin and Brown, 2016; Sheppard et al., 2019; Siraj et al., 2025). We shall discuss the implications of 2017 OF201’s orbit to Planet Nine in Section IV.3.
@@
Figure 2: Left: Plan view of the orbits of TNOs with extremely wide orbits. The newly discovered 2017 OF201, highlighted in red, is an outlier to the apsidal clustering of the others. For reference, the most probable orbit of Planet 9 / Planet X from Siraj et al. (2025) is shown in black. Right: Distribution of longitude of perihelion $\varpi$ of TNOs from Siraj et al. (2025). The purple shaded region indicates the clustering of $\varpi$ around 60. Objects with (meta)stable orbits are shown as solid (open) symbols.
III.2 Color and variability
@@
2017 OF201 has a color index of ggrr = $7±\pm0$.08 mag, which is much redder than the Sun (0.44; Willmer, 2018) and is consistent with other TNOs with scattering-disk and detached orbits (Sheppard, 2010; Ofek, 2012). Though it is on the red side of the distribution, it is not yet as red as Sedna, which has BBVV = 1.24. Given the precision and time sampling of the DECam and CFHT data, we do not detect any variability over 0.1 mag, indicating that the shape of 2017 OF201 is likely close to spherical. Deeper imaging with higher cadence is still needed to probe the short-time variability due to rotation or satellite transit.
III.3 Size and mass estimate
@@
Future observations of thermal emission from ALMA will allow for simultaneous determination of size and albedo without the need to rely on the assumed correlation (e.g., Gerdes et al., 2017). Occultation is another way to directly measure size (e.g., Rommel et al., 2020), but the next possible occultation event for stars brighter than 21 mag will be in 4 years (Nov 27 2029, 19-mag star), according to the current orbital fitting.
The discovery of such a large object with high eccentricity has interesting implications for the population of objects in trans-Neptunian space. Throughout the orbit, 2017 OF201 is only close enough to be detectable in DECaLS for $\sim$0.5% of the period, which suggests that a substantial population of about 200 similar objects – with large sizes, wide orbits, and high eccentricities – should exist but are just too difficult to detect due to their large distance.
[…]
Figure 3: Left: TNOs with wide orbits (a>a>150 au). 2017 OF201 is the second brightest in this population and has the widest orbit among all dwarf planet candidates. Right: Schematic of a possible migration pathway for 2017 OF201. The object may first have been scattered to a large $a$ while maintaining a low $q$ and then galactic tidal torques slowly increase $q$, detaching the orbit from Neptune.
IV Orbital dynamics
IV.1 Possible origin
The semi-major axis of 2017 OF201 evolve through chaotic diffusion driven by interactions with Neptune. However, the diffusion timescale $T$ increases steeply with perihelion distance. Following Hadden and Tremaine (2024), it is approximately